OpenStudy (anonymous):

a certain mountain has an elevation of 19,124 feet. in 1922, the glacier on this peak covered 6 acres. by 2000 this glacier had melted to only 1 acre. A) assume that this glacier melted at a constant rate each year, find this yearly rate. B) use your answer from part (a) to write a linear equation that gives the acreage A of this glacier t years past 1922.

OpenStudy (anonymous):

USE THIS TO WORK UR PROBLEM OUT Number of years=2002-1913=89 amount of glacier melted during 89 years = 4 -1=3 acres the yearly rate of change in acres/year =3/89 =0.034 a linear equation that gives the acreage a of this glacier t years past 1913 = 4 - 0.034 t

OpenStudy (anonymous):

thank you